OpenStudy (anonymous):

Hi, I need help with this problem...like ASAP. If anyone can help me...please :-) Here it is: How many gallons of a 3% salt solution must be mixed with 50 gallons of a 7% solution to obtain a 5% solution?

OpenStudy (anonymous):

Short Answer: 50 do you want the long anwser? that explains it step by step...

OpenStudy (anonymous):

Hi Lolly, Yes please. Thank you.

OpenStudy (anonymous):

\[\frac{ 50*0.03 + x*0.07 }{ 50+x } = 0.05\]

OpenStudy (anonymous):

Thank you FredrickV

OpenStudy (anonymous):

Long Answer: We can set up an equation: .03x + .07 * 50 = .05(x+50), where x is the number of gallons of 3% salt solution to be added Multiply both sides by 100 to make things easier 3x +7*50 = 5(x+50) 3x + 350 = 5x + 250 subtract 3x from both sides 350 = 2x + 250 subtract 250 from both sides 100 = 2x divide both sides by 2 50 = x or x = 50

OpenStudy (anonymous):

Wow! Lolly, that's just what I was looking for...thank you thank you thank you.

OpenStudy (anonymous):

no prob!:)